OpenStudy (anonymous):

prove 7 divides 2n-1 = 3divides n

OpenStudy (anonymous):

Prove tht \[\forall n \in \mathbb{N} , 7|(2^n-1) <=> 3|n\]

OpenStudy (anonymous):

a hint is to look at the multiplication table (for \[\mathbb{Z}/7\mathbb{Z}\] and computer powers of 2 mod 7, and to notice how they repeat

OpenStudy (anonymous):